OpenStudy (anonymous):

Which of the following is the solution of log (x - 3)0.001 = -3 A) x = 5 B) x = 7 C) x = 13 D) x = 15

OpenStudy (anonymous):

i think you are supposed to solve \[x-3=10\]

OpenStudy (anonymous):

since \(\log_{10}(.001)=-3\)

OpenStudy (anonymous):

on the other hand, it seems like you can always pick C as an answer, because it is C maybe 94% of the time

OpenStudy (anonymous):

\[0.001\log(x-3) = -3\] \[\log(x-3) = -3000\] \[x-3 = \frac{ 1 }{ e ^{3000} }\] \[x = 3 + \frac{ 1 }{ e^{3000} }\]

OpenStudy (anonymous):

Or wait, I'm not sure what you're really asking...go with sat.

OpenStudy (anonymous):

I think x-3 is the base of the log, \[\log_{x-3} 0.001 = 3\]Which you can rearrange as follows: \[(x-3)^{-3}=0.001\]And solve to get x=13 as mentioned

OpenStudy (anonymous):

Ah, yes that makes much more sense :)

OpenStudy (isaiah.feynman):

Clarity of questions is essential.